5. A cubical box has each edge 10 cm and another cu'boidal box is 12.5 cm wng, 10 cm wU:ie and 8 cm high.
(i) Which box has the greater lateral surface area and by
how much?
(ii) Which box has the smaJJer total surface area and by
how much?

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
SoL For cubical box: Lateral surface = 4(10>2 cm2 = 400 cm.2 •••(a) 1btal surface area = 6(10)2 cm.2 = 600 cm.2 •••(c) For cuboidal box: Lateral surface = 2(12.5 + 10) x 8 cm2 = 360 cm.2 ...(b) 'lbtal surface area = 2(12.5 x 10 + 10 x 8 + 12.5 x 8) cm.2 = 610 cm.2 •••(d) (i) Cubical box bas greater lateral surface area [From (